23. .Everyday Beryl’s produce 80 unit chocolate in month with amount RM 900. That means Beryl's can sold RM 960.The fixed operating cost is RM 100,00. How many units need to be sold to break even?



1
Expert's answer
2021-03-31T07:34:16-0400
"\frac{960-900}{80}=0.75"

"\frac {100}{0.75}=133"